# MOWA: Multiple-in-One Image Warping Model

### v1

@@ 3.5 Training Loss
4 Experiments
@@
To demonstrate the effectiveness of the proposed multiple-in-one image warping method, we evaluate its performance on six representative distorted types, including stitched images, rectified wide-angle images, unrolling shutter images, rotated images, fisheye images, and portrait photos.
4.1 Experimental Settings
@@
Implementation Details. We train the proposed model using the Adam optimizer with the momentum terms of $(0.9,0.999)$ on 8 NVIDIA A100 GPUs. The learning rate starts with a linear warm-up in the first three epochs and then decays from $1e^{-4}$ to $1e^{-6}$ following a cosine schedule in the remaining epochs. The batch size is set as 64. For the network architecture, the Transformer blocks with shifted windows [22, 32] are used in both the encoder and decoder except for the input projection layer and output projection layer. In TPS prediction heads, we adopt the convolution layers with different kernels to predict increasing numbers of control points, and the numbers are set to $10\times 10$, $12\times 12$, $14\times 14$, and $16\times 16$. More details of our network can be found in supplementary material. The complete framework is trained with a fixed input size of $256\times 256$. At the first 10 epochs, we solely train and supervise the TPS prediction heads with the point-based task classifier. Afterwards, all modules are trained collectively. During inference, MOWA supports the image warping for any resolution by scaling the predicted TPS control points and residual flow.
Metrics. Following previous works, we select PSNR and SSIM as metrics to quantitatively measure the quality of the warped results. In particular, for the portrait photos, the ShapeAcc metric is applied as suggested in Tan et al. [30]. It is specially designed for the quality of face correction, which calculates the similarity between corrected portraits and the stereographic projection of its original input.
4.2 Comparison Results
We compare the proposed MOWA with recent SotA methods on each task, including Deep_Rect [23], He et al. [12], RecRec [18], Feng et al. [7], Shih et al. [29], Tan et al. [30], and Zhu et al. [40].
Qualitative Comparison. As shown in Fig. 4, we visualize the comparison results of different methods on the testing datasets. These qualitative results demonstrate that our multiple-in-one method can handle well across various tasks, scenes, and resolutions, compared with the SotA methods specially designed for each single task. For example, for the rotated images, our method can rearrange the input to a rectangle one while keeping the original geometric layout reasonable. On the contrary, distorted buildings can be observed in the results of previous works [12, 23], in which the physical world rules such as the horizon are also perturbed. For other rectangling tasks like the stitched image, unrolling shutter image, and rectified wide-angle image, our method shows a better visual appearance especially in the image boundaries, allowing promising structural integrity among the comparison methods. In some challenging cases such as the first and second rows in stitched images, the image boundaries are dramatically stretched, but our method can still warp the images to expected structures. One important reason is that MOWA learns the generalized warping strategy from different tasks since it can extract some common knowledge from them. In addition, our MOWA mitigates the difficulty of motion estimation by disentangling it at both the region level and pixel level. Consequently, diverse structures of motions can be progressively approximated and the image details can be preserved. For the fisheye image and portrait photos, MOWA is capable of recovering the realistic geometric distribution from the inputs, despite the radial distortion or perspective distortion.
Table 1: Quantitative evaluation of the proposed multiple-in-one framework MOWA to the SotA image warping models.
Warping Tasks Metrics
Stitched Image Deep_Rect [23] 21.28 0.7140 - 52.14M
Ours 20.42 0.6307 - 49.93M
Rectified Wide-Angle Image RecRecNet [18] 18.68 0.5450 - 62.70M
Ours 18.78 0.5479 - 49.93M
Unrolling Shutter Image RecRecNet [18] 21.48 0.7602 - 62.70M
Ours 21.83 0.7798 - 49.93M
Rotated Image He et al. [12] 17.63 0.4880 - -
Deep_Rect [23] 19.89 0.5500 - 52.14M
Ours 20.83 0.5874 - 49.93M
Fisheye Image Feng et al. [7] 21.72 0.7167 - 11.65M
Ours 23.92 0.8076 - 49.93M
Portrait Photos Shih et al. [29] - - 97.253 -
Tan et al. [30] - - 97.490 -
Zhu et al. [40] - - 97.491 8.79M
Ours - - 97.475 49.93M
Quantitative Comparison. We report the quantitative evaluation results in Table 1. The proposed multiple-in-one image warping jointly learns six tasks and achieves promising performance compared with the single-task methods. For example, MOWA outperforms the SotA methods in rectified wide-angle images, unrolling shutter images, rotated images, and fisheye images. Although the purposes and motion structures of these tasks differ from each other, the proposed method can enable a unified image warping framework with an elaborately designed hierarchical motion estimation architecture and task-aware prompt learning strategy. Moreover, MOWA achieves comparable image warping performance for the stitched image and portrait photos, without intolerable performance degradation when involving more tasks and data. However, previous methods [18, 12, 7, 23, 30] tailor the specific knowledge into their models to address the single image warping task, lacking the generalization and flexibility to real-world applications.
Computation Complexity Comparison. In Table 1, we also compare the computation complexity of previous methods which have available models, and our MOWA. The results demonstrate our model size is reasonable as a multiple-in-one image warping framework. Even compared to the SotA models designed for the specific task [23, 18], our MOWA has fewer parameters to achieve better or comparable warping performance. The underlying reason is that the proposed motion estimation module discards the heavy fully connected layers and replaces them with convolutional layers. Then, the predicted region-level TPS points are further compensated with the pixel-level displacement from a compact convolutional decoder. Besides, the shared knowledge across different tasks can also help to decrease the parameter requirements of a multi-task model.
4.3 Ablation Study
Considering the aim of a multiple-in-one framework is to achieve holistic performance across various tasks, we mainly compare the different variants of the framework in terms of the average warping metrics. Additionally, the same image quality metrics (PSNR and SSIM) are shared in the first five tasks, but the portrait correction task has its own metrics like ShapeAcc. Thus, the average PSNR and SSIM from the first five tasks are reported in this part.
Table 2: Ablation study on the proposed motion estimation module. The baseline represents the predicted control points with a size of $12\times 12$. “10x10x10x10” means 4 heads are applied and each head predicts the control points with a size of $10\times 10$. Other settings are also presented in this form. “Ours” denotes the combination of “10x12x14x16” and residual flow.
Metrics Baseline 12x12x12x12 14x14x14x14 16x16x16x16 10x12x14x16 Ours
PSNR 20.29 20.38 20.42 20.02 20.48 20.84
SSIM 0.6279 0.6311 0.6406 0.6148 0.6418 0.6572
Motion Estimation. It is challenging to estimate multiple motions in one model since the motion’s complexities and patterns significantly differ across various tasks. For this purpose, we propose a flexible and hierarchical architecture to disentangle the motion estimation at the region level and pixel level. As shown in Fig. 5, better localization performance of the formed mesh can be achieved by increasing the number of TPS points. Besides, the pixel-level residual flow can provide a higher degree of freedom for the motion than only the region-level motion representation, improving the warping results, particularly in the image boundaries and details. Table 2 quantitatively demonstrates the effectiveness of the proposed hierarchical motion estimation module. We also found the upper bound occurs when increasing the number of control points, which suggests the performance of multiple-in-one image warping would be limited without proper decoupling motion manner.
Table 3: Ablation study on different task classifiers for the multiple-in-one image warping framework.
Metrics w/o Classifier Image-based Classifier Point-based Classifier Ours
PSNR 20.48 20.58 20.63 20.83
SSIM 0.6418 0.6451 0.6463 0.6558
Parameters - 3.39M 0.0592M 0.0594M
Task Classifier. As the image-based classifier involving redundant image features burdens the main image warping framework, we propose a lightweight point-based classifier to learn the task type from each input image. As listed in Table 3, four baseline models are designed: image warping model without the task classifier, with the image-based classifier, with the point-based classifier, and with the point-based classifier compensated with the pooling global features (Ours). Note that we validate different task classifiers only with the TPS prediction modules, showing their direct influence on the region-level motion estimation. The quantitative results demonstrate that we can obtain evident performance gain beyond the vanilla network by adding the task classifier, indicating that the information on task type is meaningful in multi-task training.
Prompt Learning. In our MOWA, a prompt learning module is designed to modulate the feature map with the soft task label, which helps to dynamically navigate its extensive parameter space to achieve task-aware image warping. By combining this module into MOWA, the averaged PSNR metric of all warping results gains $+0.21$dB improvements beyond the baseline. More experimental results can be found in supplementary materials.
Figure 6: Generalization evaluation of the proposed MOWA: cross-domain evaluation (top) and zero-shot evaluation (bottom). The visualized flow and control points are both predicted by MOWA. In image retargeting results, the red dotted lines measure the stretch extent of face or body by warping operations.
4.4 Generalization
We show the generalization ability of the proposed MOWA in terms of the cross-domain and zero-shot evaluations in Fig. 6. For the cross-domain evaluation, the new inputs belong to the above six practical image warping tasks but they are captured in real-world settings with different cameras, resolutions (can be up to 4K), and scenes. For the zero-shot evaluation, we consider a new image warping task, i.e., image retargeting, which aims to flexibly change the image scale without distorting the content as much as possible.
The visualization results demonstrate that MOWA can well extend to real-world scenarios though its training datasets are mostly synthesized by hand-crafted camera models or manipulation spaces. One possible reason is the multiple-in-one model can naturally address the overfitting issue on specific datasets by learning various tasks. In addition to the cross-domain evaluations, we find that while our model does not involve the image retargeting task during training, it is still able to warp the image based on the “content-aware” principle. As we can observe, the predicted control points are accurately aligned to the image boundaries. Such knowledge transferring to new tasks potentially benefits from the shared motion perception across different tasks. Therefore, our results show fewer geometrical distortions for the foreground than the crop and resize operation, with the face and body experiencing less stretching.
5 Conclusion
We have proposed MOWA in this work, the first multiple-in-one image warping framework in the field of computational photography. It considers six representative and practical tasks in one learning model and uses a unified motion representation to achieve various warping purposes. In particular, to mitigate the difficulty of approximating diverse motions of different tasks, we propose to disentangle the motion estimation at both the region level and pixel level. Then, we enable MOWA the explicit task-aware ability by introducing a lightweight point-based classifier. Compared to the common image-based classifier, it can achieve comparable performance while offering significant parameter reduction. Subsequently, we feed the soft task label predicted by the task classifier into a prompt learning module and further modulate the feature maps in the decoder, which facilitates a single model to efficiently navigate and leverage its extensive parameter space to meet various warping requirements. Comprehensive experiments demonstrate that MOWA outperforms different SotA methods specifically designed for each single task, with an affordable model size.
In the future, we plan to empower MOWA with cross-view and cross-modal abilities, aiming to build a foundation model for universal image warping.
Appendix 0.A Supplementary Materials
0.A.1 Overview
In this document, we provide the following supplementary content:
  • More implementation details of the multiple-in-one framework (Section 0.A.2).
  • More results and analysis of ablation study (Section 0.A.3 and Section 0.A.4).
  • More qualitative results of the comparison methods (Section 0.A.5).
  • More generalization evaluations (Section 0.A.6).
  • Limitation discussion (Section 0.A.7).
0.A.2 More Implementation Details
@@
We design the architecture of our MOWA based on the encoder-decoder style, it enables both region-level control point regression and pixel-level residual flow prediction. To be more specific, the Transformer blocks with shifted windows (with the size of 8) are exploited in each layer except for the input and output projection layers. The basic dimension of channels is set to 32 and linearly increases along the layers in the encoder network, which is oppositely decreased to 2 in the decoder network. During training, all inputs are sized to $256\times 256$ and the size of the last feature maps in the encoder network is $16\times 16$. Moreover, the depths of each Transformer block are set to 2 and the head numbers of multi-head self-attention are $[1,2,4,8,16,16,8,4,2]$ along the whole layers. To establish a compact network, the convolutional layers are used to regress the coordinates of 2D control points, which enable significant parameter reduction compared with the fully connected layers. The configuration details of these regression heads are listed in Table 4. For the lightweight point-based classifier, three 1D convolutional layers with channel dimensions of $256,256,512$ are used to extract the features of input points and then three fully connected layers with unit numbers of $512,256,6$ are used to classify their task types.
In this work, our primary objective is to contribute the first multiple-in-one image warping model to the computational photography community. The aforementioned configuration has been proven effective in simultaneously learning six tasks within a single model. It achieves superior or comparable performance to the state-of-the-art single-task methods even with fewer parameters. We believe more complex architectures could further improve the image warping performance and leave it as one of the future works.
@@
Table 4: The configurations of convolutional layers to regress different sizes of control points (the size of input feature maps is $16\times 16$).
@@
0.A.3 More Analysis and Ablations on Multi-Task Learning
@@
To intuitively analyze the influence of multi-task learning and the effectiveness of the proposed prompt learning, we visualize the quantitative metrics of the warping results of validation images for each task. As illustrated in Figure 7, the values of normalized PSNR and SSIM are plotted along different training epochs. Considering the ranges of PSNR and SSIM of different warping tasks significantly differ from each other, we normalize all values into the range of $[0,1]$ to eliminate the data bias. For the portrait photos, we leverage the available warping flow in its training dataset and obtain the corrected images to compute the PSNR and SSIM. For other tasks, we directly use the ground truth of warped images in the corresponding datasets.
From Figure 7 (a), we have the following three observations: (1) Different tasks show various levels of difficulty when training a multiple-in-one image warping model. For example, learning to warp the unrolling shutter image (task 3) is generally easier than other tasks, which shows the fastest convergence at the first 80 epochs. The reason is that the structures of unrolling shutter images are basically regular, where more than two boundaries are straight and do not need to be rectified. On the contrary, learning to warp the stitched images (task 1), rectified wide-angle images (task 2), and portrait photos (task 6) is more challenging (slow convergences can be observed) since their boundaries or warping motions vary greatly in datasets. Especially in the portrait photo dataset, the numbers, shapes, and locations of the faces are quite diverse. (2) The multiple-in-one model tends to sacrifice the performance of some individual tasks to achieve the improvement of holistic warping performance. Particularly, the accuracy of warping the unrolling shutter image is dramatically reduced to improve the warping performance of other tasks since the $80$th epoch. Such a trade-off among different tasks facilitates an overall improvement but leads to performance conflicts between some tasks. (3) The relationship of different tasks can be positive or negative. For example, the curves of the stitched image and rotated image show consistent trends during the MOWA’s training as they share the similar warping principle, i.e., rectangling the irregular image boundaries and keeping the content unchanged. Thus, meaningful interactions between these two tasks could happen if learning a multiple-in-one model. However, the curve of warping the fisheye image shows a converse trend to those of the stitched image and rotated image. The reason for this is that correcting the radial distortion present in fisheye images significantly alters the scene’s layout while leaving the image boundaries almost untouched. This approach contrasts with the foundational principles of tasks based on image rectangling. As a result, the imbalanced convergences can be noticed across these tasks with negative relationships.
Figure 8: t-SNE visualization of the learned prompts in MOWA.
To dynamically boost the task-aware image warping in a single model, we propose a prompt learning module and a point-based task classifier. As shown in Figure 7 (b), the performance conflicts of different tasks are relieved by prompt learning. All tasks show a similar improvement trend as the training epoch increases, without dramatic performance degradation in certain individual tasks. More importantly, the multiple-in-one framework achieves the unified and best warping performance on all tasks at the end of training epochs. This phenomenon suggests that the framework knows to discriminate and warp different input types using the learned task-specific prior knowledge. Our designed prompts enable MOWA to efficiently traverse and harness its vast parameter space to meet various warping requirements. We also visualize the t-SNE of the learned prompts of MOWA in Figure 8. We can observe these prompts are well-clustered based on the task types, demonstrating the ability of prompts to learn discriminative motion context that aids in holistic image warping.
Figure 9: Ablation study on TPS regression heads. We show different warping performances from two groups of control points. For each comparison, we show the warping results from the control points with the size of $16\times 16$, $16\times 16$, $16\times 16$,16×$\times 16$ 16 (top), and $10\times 10$, $12\times 12$, $14\times 14$, $16\times 16$ (bottom). The arrows highlight the under-rectified distortion parts.
0.A.4 More Analysis and Ablations on Motion Estimation
To eliminate the difficulty of estimating various warping motions, we propose to disentangle the motion estimation at both the region level and pixel level. In the region-level estimation, the hierarchical TPS control points with increasing numbers are designed to gradually approximate the motion structure. As shown in Figure 9, the results of the ablation study demonstrated the proposed TPS regression heads achieve more accurate warping performance on severe geometric distortion correction, compared with the settings with even more control points. This indicates the upper bound of the control point numbers exists in certain warping tasks, where adding more control points would restrict the model’s flexibility to estimate complex motions. The MOWA, by contrast, learns the progressive warping rules in a spatially coarse-to-fine manner, allowing more flexible structures at the beginning levels of cascade regression heads.
0.A.5 More Comparison Results
We show more qualitative comparison results in Figure 10, Figure 11, Figure 12, and Figure 13. The results demonstrate the proposed MOWA shows superior performance across diverse tasks, scenes, and geometric distortions, verifying its promising practicality beyond previous methods customized to a single task.
Figure 10: More comparison results on the stitched image test dataset.
Figure 11: More comparison results on the rotated image test dataset.
Figure 12: More comparison results on the rectified wide-angle image, unrolling shutter image, and fisheye image test datasets. For each comparison, we show the input image, results of comparison methods [18, 7], MOWA, and ground truth from left to right.
Figure 13: More comparison results on the portrait photo test dataset.
0.A.6 More Generalization Evaluations
In addition to the performance on the test datasets, we demonstrate the proposed MOWA has favorable generalization ability to unfamiliar domains and unseen tasks. In particular, we show cross-domain evaluations of MOWA in Figure 14, Figure 15, Figure 16, and Figure 17. The cross-domain datasets are derived from real-world datasets [25, 18, 35, 40]. Compared to the training datasets, the images in real-world datasets are captured by various cameras or obtained by practical manipulations. As shown in these figures, while the significant domain gap exists, the proposed approach can well extend to other domains. It is noticed that MOWA also exhibits satisfied robustness to the noisy data. For instance, in Figure 17, although MOWA is trained on the synthetic dataset with clean masks, the warping results of the real-world dataset with noisy pseudo masks (the clean masks are not available in some practical applications) are still structurally reasonable.
The zero-shot evaluation is shown in Figure 19. We consider a new image retargeting task as an unseen domain to validate the MOWA’s generalization. Compared to the resize operation, MOWA can retarget the resolution of the input image with less stretching of persons and objects. Compared to the classical image rectangling method [12], fewer geometric distortions are introduced by warping the input image with MOWA.
Figure 14: Cross-domain evaluation on the real-world high-resolution stitched images. The resolution of the input images ranges from $1191\times 722$ to $3616\times 2048$.
Figure 15: Cross-domain evaluation on the real-world fisheye images.
Figure 16: Cross-domain evaluation on the real-world portrait photos.
@@ 0.A.6 More Generalization Evaluations
Figure 17: While MOWA is trained on the synthetic datasets with clean masks, it shows good robustness to the real-world data with noisy pseudo masks. The motion estimation of real-world images, e.g., the predicted control points, and warping flow, exhibit structurally reasonable distributions similar to those of the synthetic images.

### v2

@@ III-E Training Loss
In summary, the proposed multiple-in-one image warping framework bring the following benefits.
  • Unlike previous task-specific image warping models, our method can recover various geometrically distortion images within a single network. It does not require prior knowledge of the camera models or manipulation spaces; it is also friendly to use and relies only on the observed input image to perform the customized image warping.
  • Our method provides greater flexibility and cost-effectiveness in real-world scenarios, unlike previous methods that need a proportionally larger model size as the number of warping tasks increases.
  • Thanks to multi-task learning, our method develops a generalized motion representation across various image warping tasks, demonstrating remarkable performance in cross-domain evaluations and unseen tasks.
IV Experiments
@@
To demonstrate the effectiveness of the proposed multiple-in-one image warping method, we evaluate its performance on six representative distorted types, including stitched images, rectified wide-angle images, unrolling shutter images, rotated images, fisheye images, and portrait photos, covering the mainstream practical image warping tasks.
IV-A Experimental Settings
@@
Implementation Details. We train the proposed model using the Adam optimizer with the momentum terms of $(0.9,0.999)$ on 8 NVIDIA A100 GPUs. The learning rate starts with a linear warm-up in the first three epochs and then decays from $1e^{-4}$ to $1e^{-6}$ following a cosine schedule in the remaining epochs. The batch size is set as 64. The complete framework is trained with a fixed input size of $256\times 256$. At the first 10 epochs, we solely train and supervise the TPS prediction heads with the point-based task classifier. Afterwards, all modules are trained collectively. During inference, the proposed method supports image warping for any resolution by scaling the predicted TPS control points and residual flow.
@@
Network Configuration. We design the image warping network based on the encoder-decoder architecture, enabling both region-level control point regression and pixel-level residual flow prediction. Specifically, the Transformer blocks with shifted windows [48, 49] are used in both the encoder and decoder except for the input projection layer and output projection layer. The basic dimension of channels is set to 32 and linearly increases along the layers in the encoder network, which is oppositely decreased to 2 in the decoder network. Moreover, the depths of each Transformer block are set to 2 and the head numbers of multi-head self-attention are $[1,2,4,8,16,16,8,4,2]$ along the whole layers. In TPS prediction heads, we adopt the convolution layers with different kernels to predict increasing numbers of control points, and the numbers are set to $10\times 10$, $12\times 12$, $14\times 14$, and $16\times 16$. The configuration details of these regression heads are listed in Table I. Such a design enables significant parameter reduction compared with the fully connected layers. For the lightweight point-based classifier, three 1D convolutional layers with channel dimensions of $256,256,512$ are used to extract the features of input and then three fully connected layers with unit numbers of $512,256,6$ are used to classify their task types.
@@
TABLE I: The configurations of convolutional layers to regress different sizes of control points (the size of input feature maps is $16\times 16$).
@@
TABLE II: Quantitative evaluation of the proposed multiple-in-one framework to the SotA image warping models.
Warping Tasks Metrics
Rotated Image He et al. [5] 17.63 0.4880 - -
Deep_Rect [6] 19.89 0.5500 - 52.14M
Ours 21.01 0.5961 - 49.93M
Rectified Wide-Angle Image He et al. [5] 15.36 0.4211 - -
RecRecNet [21] 18.68 0.5450 - 62.70M
Ours 18.69 0.5450 - 49.93M
Stitched Image He et al. [5] 14.70 0.3775 - -
Deep_Rect [6] 21.28 0.7140 - 52.14M
Ours 20.72 0.6425 - 49.93M
Unrolling Shutter Image RecRecNet [21] 21.48 0.7602 - 62.70M
Ours 21.69 0.7795 - 49.93M
Fisheye Image PCN [37] 21.37 0.6925 - 26.19M
Feng et al. [3] 21.72 0.7167 - 11.65M
Ours 22.25 0.7488 - 49.93M
Portrait Photos Shih et al. [41] - - 97.253 -
Tan et al. [25] - - 97.490 -
Zhu et al. [26] - - 97.491 8.79M
Ours - - 97.477 49.93M
Metrics. Following previous works, we select PSNR and SSIM as metrics to quantitatively measure the quality of the warped results. Please note that it is challenging to use the Average Endpoint Error (EPE) metric to evaluate the motion estimation performance in practical image warping tasks, because the accurate labels of motion are hard to obtain and unavailable in all the above test datasets. As a consequence, most previous methods have opted to learn the motion in an unsupervised manner and supervise the image warping model at the warped pixel level.
For the portrait correction task, the ShapeAcc metric is applied as suggested in Tan et al. [25]. It is specially designed for the quality of face correction, which calculates the similarity between corrected portraits and the stereographic projection of its original input.
IV-B Comparison Results
We compare the proposed MOWA with recent SotA methods on each task, including Deep_Rect [6], He et al. [5], RecRecNet [21], PCN [37], Feng et al. [3], Shih et al. [41], Tan et al. [25], and Zhu et al. [26].
Qualitative Comparison. As shown in Fig. 4, we visualize the comparison results of different methods on the testing datasets. These qualitative results demonstrate that our multiple-in-one method can handle various tasks, scenes, and resolutions well, compared with the SotA methods specially designed for each task. For example, for the rotated images, our method can rearrange the input to a rectangle one while keeping the original geometric layout reasonable. On the contrary, distorted buildings can be observed in the results of previous works [5, 6], in which the physical world rules such as the horizon are perturbed. For other rectangling tasks like the stitched image, unrolling shutter image, and rectified wide-angle image, our method shows a better visual appearance, especially in the image boundaries, allowing promising structural integrity among the comparison methods. In some challenging cases, such as the first and second rows in stitched images, the image boundaries are dramatically stretched, but our method can still warp the images to expected structures. One important reason is that MOWA learns the generalized warping strategy from different tasks since it can extract some common knowledge from them. In addition, our method mitigates the difficulty of motion estimation by disentangling it at both the region level and pixel level. Consequently, diverse structures of motions can be progressively approximated and the image details can be preserved. For the fisheye image and portrait photos, MOWA is capable of recovering the realistic geometric distribution from the inputs, despite the radial distortion or perspective distortion. Please refer to more visual comparison images, interactive warping visualizations, and dynamic warping results on the project page: https://kangliao929.github.io/projects/mowa/.
Quantitative Comparison. We report the quantitative evaluation results in Table II. The proposed multiple-in-one image warping jointly learns six tasks and achieves promising performance compared with the single-task methods. For example, MOWA outperforms the SotA methods in rectified wide-angle images, unrolling shutter images, rotated images, and fisheye images, thanks to the elaborately designed hierarchical motion estimation architecture and task-aware prompt learning strategy. Moreover, MOWA achieves comparable image warping performance for stitched images and portrait photos without intolerable performance degradation when involving more tasks and data. The results suggest the generalizability and flexibility of MOWA, which are not achievable by previous methods [21, 5, 3, 6, 25] that tailor the specific knowledge into their models to address the single image warping task.
Computation Complexity Comparison. In Table II, we also compare the computation complexity of the proposed method with previous methods that make their models available. The comparison suggests that our model size is reasonable and affordable as a multiple-in-one image warping framework. Even compared to the SotA models designed for the specific task [6, 21], our MOWA has fewer parameters to achieve better or comparable warping performance. The underlying reason is that the shared knowledge across different tasks can relieve the burden of the parameter requirements of a multi-task model. Besides, the proposed motion estimation module discards the heavy fully connected layers and replaces them with convolutional layers. Then, the predicted region-level TPS points are further compensated with the pixel-level displacement from a compact convolutional decoder.
IV-C Ablation Study
Considering the aim of a multiple-in-one framework is to achieve holistic performance across various tasks, we mainly compare the different variants of the framework in terms of the average warping metrics. Additionally, the same image quality metrics (PSNR and SSIM) are shared in the first five tasks, but the portrait correction task has its own metrics like ShapeAcc. Thus, the average PSNR and SSIM from the first five tasks are mainly reported in this part.
TABLE III: Ablation study on the proposed motion estimation module. The baseline represents the predicted control points with a size of $12\times 12$. “10-10-10-10” means 4 heads are applied and each head predicts the control points with a size of $10\times 10$. Other settings are also presented in this form. “Ours” denotes the combination of “10-12-14-16” and residual flow.
Metrics Baseline 12-12-12-12 14-14-14-14 16-16-16-16 10-12-14-16 Ours
PSNR 20.29 20.38 20.42 20.02 20.48 20.84
SSIM 0.6279 0.6311 0.6406 0.6148 0.6418 0.6572
Motion Estimation. It is challenging to estimate multiple motions in one model since the motion’s complexities and patterns significantly differ across various tasks. For this purpose, we proposed a flexible and hierarchical architecture to disentangle the motion estimation at the region level and pixel level. As shown in Fig. 5, better localization performance of the formed mesh can be achieved by increasing the number of TPS points. Besides, the pixel-level residual flow can provide a higher degree of freedom for the motion than only the region-level motion representation, improving the warping results, particularly in the image boundaries and details. Table III quantitatively demonstrates the effectiveness of the proposed hierarchical motion estimation module. We also found the upper bound occurs when continuously increasing the number of control points, e.g, the performance of four motion estimation heads to predict the size of $16\times 16$ TPS even worse than the size of $12\times 12$ TPS. This suggests the performance of multiple-in-one image warping would be limited without proper decoupling motion manner.
TABLE IV: Ablation study on different task classifiers for the multiple-in-one image warping framework.
Metrics w/o Classifier Classifier-Image Classifier-Point Ours
PSNR 20.48 20.58 20.63 20.83
SSIM 0.6418 0.6451 0.6463 0.6558
Parameters - 3.39M 0.0592M 0.0594M
@@ IV-C Ablation Study
Task Classifier. As the image-based classifier involving redundant image features burdens the main image warping framework, we propose a lightweight point-based classifier to learn the task type from each input image. As listed in Table IV, four baseline models are designed: image warping model without the task classifier, with the image-based classifier, with the point-based classifier, and with the point-based classifier compensated with the pooling global features (Ours). Note that we validate different task classifiers only with the TPS prediction modules, showing their direct influence on the region-level motion estimation. The quantitative results demonstrate that we can obtain evident performance gain beyond the vanilla network by adding the task classifier, indicating that the information on task type is meaningful in multi-task training.
Figure 7: Generalization evaluation of the proposed method: cross-domain evaluation (top) and zero-shot evaluation (bottom). The visualized flow and control points are both predicted by MOWA. In image retargeting results, the red dotted lines measure the stretch extent of the face or body by warping operations.
IV-D Analysis on Multi-Task Learning and Prompt Learning
In the proposed method, a prompt learning module is designed to modulate the feature map with the soft task label, which helps to dynamically navigate its extensive parameter space to achieve task-aware image warping. By combining this module into MOWA, the averaged PSNR metric of all warping results gains $+0.21$dB improvements beyond the baseline.
@@ IV-D Analysis on Multi-Task Learning and Prompt Learning
To further analyze the influence of multi-task learning and the effectiveness of the proposed prompt learning, we visualize the quantitative metrics of the warping results of validation images for each task. As illustrated in Figure 6, the values of normalized PSNR and SSIM are plotted along different training epochs. Considering the ranges of PSNR and SSIM of different warping tasks significantly differ from each other, we normalize all values into the range of $[0,1]$ to eliminate the data bias. For the portrait photos, we leverage the available warping flow in its training dataset and obtain the corrected images to compute the PSNR and SSIM. For other tasks, we directly use the ground truth of warped images in the corresponding datasets.
@@
Figure 8: While our method is trained on the synthetic datasets with clean masks, it shows good robustness to the real-world data with noisy pseudo masks. The motion estimation of real-world images, e.g., the predicted control points, and warping flow, exhibit structurally reasonable distributions similar to those of the synthetic images.
From Figure 6 (a), we have the following three observations: (1) Different tasks show various levels of difficulty when training a multiple-in-one image warping model. For example, learning to warp the unrolling shutter image (task 3) is generally easier than other tasks, which shows the fastest convergence at the first 80 epochs. The reason is that the structures of unrolling shutter images are basically regular, where more than two boundaries are straight and do not need to be warped. On the contrary, learning to warp the stitched images (task 1), rectified wide-angle images (task 2), and portrait photos (task 6) is more challenging (slow convergences can be observed) since their boundaries or expected motions vary greatly in datasets. Especially in the portrait photo dataset, the numbers, shapes, and locations of the faces are quite diverse. (2) The multiple-in-one model tends to sacrifice the performance of some individual tasks to achieve the improvement of holistic warping performance. Particularly, the accuracy of warping the unrolling shutter image is dramatically reduced since the 80th epoch, but the performance of other tasks continues to improve. Such a trade-off among different tasks facilitates an overall improvement but leads to performance conflicts between some tasks. (3) The relationship of different tasks can be positive or negative. For example, the curves of the stitched image and rotated image show consistent trends during the MOWA’s training as they share the similar warping principle, i.e., rectangling the irregular image boundaries and keeping the content unchanged. Thus, meaningful interactions between these two tasks could happen if learning a multiple-in-one model. However, the curve of warping the fisheye image shows a converse trend to those of the stitched image and rotated image. The reason for this is that correcting the radial distortion present in fisheye images significantly alters the scene’s layout while leaving the image boundaries almost untouched. This approach contrasts with the foundational principles of tasks based on image rectangling. As a result, the imbalanced convergences can be noticed across these tasks with negative relationships.
To dynamically boost the task-aware image warping in a single model, we propose a prompt learning module and a point-based task classifier. As shown in Figure 6 (b), the performance conflicts of different tasks are relieved by prompt learning. All tasks show a similar improvement trend as the training epoch increases, without dramatic performance degradation in certain individual tasks. More importantly, the multiple-in-one framework achieves the unified and best warping performance on all tasks at the end of training epochs. This phenomenon suggests that the framework knows to discriminate and warp different input types using the learned task-specific prior knowledge. Our designed prompts enable MOWA to efficiently traverse and harness its vast parameter space to meet various warping requirements.
We also visualize the t-SNE of the learned prompts of MOWA in Figure 6 (c). We can observe these prompts are well-clustered according to the task types. This clear clustering demonstrates the ability of the prompts to learn and represent discriminative motion context, which significantly aids in holistic image warping. The visualization underscores the effectiveness of our approach in capturing and leveraging task-specific features to enhance model performance.